OpenStudy (anonymous):

Find the distance between the points given. (0, -6) and (9, 6) 9 15 √(26)

OpenStudy (anonymous):

9-0=9 6-(-6)=12 a^2+b^2=c^2 9^2+12^2=c^2 81+144=c^2 225=c^2 \[\sqrt{225}=c\] c=15